# SmartThinker: Learning to Compress and Preserve Reasoning by Step-Level Length Control

### v1

@@ Introduction
To address this issue, previous studies have attempted to penalize the overall length of reasoning chains within a RL framework to encourage the generation of more concise reasoning chains (Aggarwal and Welleck 2025; Arora and Zanette 2025; Hou et al. 2025; Zhang and Zuo 2025). However, these approaches typically focus only on the overall length of the final output and lack fine-grained feedback on intermediate reasoning steps. Under such a sparse reward signals, models are prone to reward hacking, where the model compresses critical reasoning steps too aggressively while retaining unnecessary details in simpler steps, ultimately compromising both reasoning efficiency and quality. As shown in Fig. 1(a), we visualize the relationship between the importance and length of each reasoning step to better understand the impact of this overall length penalty mechanism. The results reveal that regardless of whether the problem is simple or complex, both the number and total length of high-importance steps remain quite limited, displaying a sharply skewed distribution. These findings indicate that the global length penalty mechanism compresses step lengths indiscriminately and fails to improve the quality of model reasoning effectively. This observation raises a key question: Can we apply differentiated length control at the step level to achieve a better balance between reasoning performance and computational efficiency?
@@
We use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B as base models and train these models on a randomly sampled subset of 3000 QA pairs from the DeepScaleR-Preview dataset (Luo et al. 2025b). For evaluation, we use test sets from six diverse reasoning benchmarks: AIME24, AIME25 (Veeraboina 2023), AMC23, MATH (Lewkowycz et al. 2022), MinervaMATH (Hendrycks et al. 2021), and Olympiad-Bench (Gao et al. 2024). Please see Appendix A.1 for detailed descriptions of these benchmarks.
@@
The hyperparameters are configured as $k_{0}=0.6$, $\gamma=0.95$, $\delta_{1}=0.03$, and $\delta_{2}=0.08$. Due to space limitation, we comprehensively introduce our experimental setting in Appendix A.4.
@@
As shown in Table 1, SmartThinker achieves the highest reasoning accuracy while using the fewest tokens. For example, compared to Alpha-0.1, SmartThinker improves the upper bound of reasoning performance on the challenging AIME24 dataset, increasing Pass@10 accuracy by 3.3% while reducing average token usage by about 1.5k. On AIME25, SmartThinker also increases Maj@10 accuracy by 3.3%, reflecting a more consistent ability to produce correct answers. These improvements stem from SmartThinker’s fine-grained control over the reasoning process, enabling a better balance between performance and computational efficiency. Notably, L1-Max, despite being trained on the more powerful DeepScaleR (Luo et al. 2025b) model, achieves only 50% Pass@10 and 30% Maj@10 on AIME2024, suggesting that its global length penalty may excessively compress critical reasoning steps, sacrificing accuracy for token savings. Moreover, Thinkless has limited improvements in reasoning accuracy on simpler datasets like AMC23 and MATH, possibly because switching to shorter reasoning hampers sufficient exploration of correct answers. In summary, these results confirm the superiority of SmartThinker, which consistently delivers more accurate reasoning with greater computational efficiency. Please refer to Appendix A.5 for the experimental results of the 7B model.
@@ A.1 Benchmark Datasets
In this section, we will describe in detail all the datasets used in our experiment.
@@ A.5 Experimental results of the 7B model
A.6 Case Studies

### v2

@@ Introduction
To address this issue, previous studies have attempted to penalize the overall length of reasoning chains within a RL framework to encourage the generation of more concise reasoning chains (Aggarwal and Welleck 2025; Arora and Zanette 2025; Hou et al. 2025; Zhang and Zuo 2025). However, these approaches typically focus only on the overall length of the final output and lack fine-grained feedback on intermediate reasoning steps. Under such a sparse reward signals, models are prone to reward hacking, where the model compresses critical reasoning steps too aggressively while retaining unnecessary details in simpler steps, ultimately compromising both reasoning efficiency and quality. As shown in Fig. 1(a), we visualize the relationship between the importance and length of each reasoning step to better understand the impact of this overall length penalty mechanism. The results reveal that regardless of whether the problem is simple or complex, both the number and total length of high-importance steps remain quite limited, displaying a sharply skewed distribution. These findings indicate that the global length penalty mechanism compresses step lengths indiscriminately and fails to improve the quality of model reasoning effectively. Please refer to Appendix A.8 for a detailed mathematical proof. This observation raises a key question: Can we apply differentiated length control at the step level to achieve a better balance between reasoning performance and computational efficiency?
@@
We use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B as base models and train these models on a randomly sampled subset of 3000 QA pairs from the DeepScaleR-Preview dataset (Luo et al. 2025b). For evaluation, we use six in-domain reasoning benchmarks: AIME24, AIME25 (Veeraboina 2023), AMC23, MATH (Lewkowycz et al. 2022), MinervaMATH (Hendrycks et al. 2021), and Olympiad-Bench (He et al. 2024) and three out-of-domain benchmarks: TruthfulQA (Lin, Hilton, and Evans 2021), RACE (Lai et al. 2017), Live-Code-Bench (Jain et al. 2024). Please see Appendix A.1 for more details.
@@
The hyperparameters are configured as $k_{0}=0.6$, $\gamma=0.95$, $\delta_{1}=0.03$, and $\delta_{2}=0.08$.Due to space limitations, detailed experimental settings are provided in Appendix A.4.
@@
As shown in Table 1, SmartThinker achieves the highest reasoning accuracy while using the fewest tokens. For example, compared to Alpha-0.1, SmartThinker improves the upper bound of reasoning performance on the challenging AIME24 dataset, increasing Pass@10 accuracy by 3.3% while reducing average token usage by about 1.5k. On AIME25, SmartThinker also increases Maj@10 accuracy by 3.3%, reflecting a more consistent ability to produce correct answers. These improvements stem from SmartThinker’s fine-grained control over the reasoning process, enabling a better balance between performance and computational efficiency. Notably, L1-Max, despite being trained on the more powerful DeepScaleR (Luo et al. 2025b) model, achieves only 50% Pass@10 and 30% Maj@10 on AIME2024, suggesting that its global length penalty may excessively compress critical reasoning steps, sacrificing accuracy for token savings. Moreover, Thinkless has limited improvements in reasoning accuracy on simpler datasets like AMC23 and MATH, possibly because switching to shorter reasoning hampers sufficient exploration of correct answers. In summary, these results confirm the superiority of SmartThinker, which consistently delivers more accurate reasoning with greater computational efficiency. See Appendices A.5 and A.6 for the 7B model and out-of-domain results.
@@ A.1 Benchmark Datasets
In this section, we will describe in detail all the datasets used in our experiment. These datasets are divided into in-domain datasets and out-of-domain datasets.
In-domain datasets are as follows:
@@
Out-of-domain datasets are as follows:
  • TruthfulQA is a benchmark to measure whether a language model is truthful in generating answers to questions. The benchmark comprises 817 questions that span 38 categories, including health, law, finance and politics. Questions are crafted so that some humans would answer falsely due to a false belief or misconception. To perform well, models must avoid generating false answers learned from imitating human texts.
  • RACE is a large-scale reading comprehension dataset with more than 28,000 passages and nearly 100,000 questions. The dataset is collected from English examinations in China, which are designed for middle school and high school students. The dataset can be served as the training and test sets for machine comprehension.
  • Live-Code-Bench problems are collected from competition programming websites with particular focus on maintaining problem quality, test case quality, and problem difficulty diversity. This scenario currently hosts over 500 problems from LeetCode, AtCoder, and Codeforces. Each problem instance consists of a problem description, input/output examples, and hidden test cases. Additionally, every problem is tagged with its difficulty level and release date, which allows measuring model performance across different time windows. The goal is to generate a correct and efficient solution for each problem instance. We randomly selected 100 samples from the dataset for quick validation.
@@ A.4 Experimental Setting Details
(a) Comparison of generalization between SmartThinker and the original model at a scale of 1.5B
(b) Comparison of generalization between SmartThinker and the original model at a scale of 7B
Figure 7: Out-of-domain generalization of SmartThinker.
@@ A.5 Experimental results of the 7B model
A.6 Out-of-Domain Generalization
Fig. 7 illustrates the generalization ability of SmartThinker across various domains, including factual question answering (TruthfulQA), reading comprehension (RACE), and code generation (Live-Code-Bench). As shown in the figure, SmartThinker performs on par with the original model on TruthfulQA, while achieving slight improvements in reading comprehension and code generation tasks. These results indicate that our proposed two-stage learnable framework not only preserves the model’s inherent knowledge, but also enhances its reasoning capability through RL in the mathematical domain. Consequently, SmartThinker becomes more adept at handling tasks that require deep rasoning. This further demonstrates that SmartThinker improves not only mathematical reasoning but also the overall performance of the model on general reasoning tasks.
A.7 Case Studies
@@
A.8 Theoretical Analysis
In this section, we provide a theoretical proof that step-level length control can achieve a better trade-off between reasoning quality and computational efficiency compared to global length control.
Typically, the step-level length control algorithm aims to maximize the following reward:
$R_{\text{step}}=\mathbb{1}(\text{correct})-\sum_{i=1}^{n}\lambda_{i}\frac{l_{i}}{T},$ (19)
where $\mathbb{1}(\cdot)$ denotes the indicator function, $\lambda_{i}$ is the length penalty coefficient for the $i$-th step, $l_{i}$ is the length of the $i$-th step, and $T=\sum_{i=1}^{n}l_{i}$ denotes the total length of the response.
The reward for the $i$-th step can therefore be defined as:
$r_{i}^{\text{step}}=m_{i}-\lambda_{i}\frac{l_{i}}{T},$ (20)
where $m_{i}$ represents the marginal contribution of the $i$-th step to the overall correctness of the answer.
In the GRPO algorithm, the advantage for the $i$-th step is defined as:
$A_{i}=\frac{r_{i}-\overline{r}}{\sigma_{r}},$ (21)
where $\overline{r}$ and $\sigma_{r}$ denote the mean and standard deviation of the rewards, respectively.
Let the current length of the $i$-th step be $l_{i}$. After an update to the policy gradient, it becomes $l_{i}^{\text{step}}$. For simplicity, we assume that the change in step length is proportional to the policy gradient in reinforcement learning, yielding:
$\Delta l_{i}\propto\mathbb{E}\left[\nabla_{\theta}\frac{\pi_{\theta}(s_{i}\mid q)}{\pi_{\text{old}}(s_{i}\mid q)}A_{i}\right].$ (22)
Considering only the variability introduced by advantage estimation, we have:
$l_{i}^{\text{step}}=l_{i}+\eta\cdot A_{i}.$ (23)
Substituting Eq. 20 and 21 into Eq. 23:
$l_{i}^{\text{step}}=l_{i}\cdot\left(1-\frac{\alpha\lambda_{i}}{T}\right)+\alpha\cdot\delta_{i},$ (24)
where $\alpha=\frac{\eta}{\sigma_{r}}$ and $\delta_{i}=m_{i}-\overline{r}$.
If we focus solely on differences in penalty coefficients, the Eq. 24 can be approximated as:
$l_{i}^{\text{step}}\approx l_{i}\cdot\left(1-\frac{\alpha\lambda_{i}}{T}\right).$ (25)
Similarly, for the global length control algorithm, after one update, the length of the $i$-th step can be expressed as:
$l_{i}^{\text{global}}\approx l_{i}\cdot\left(1-\frac{\alpha\lambda}{T}\right).$ (26)
Let $v_{i}$ denote the contribution density of the $i$-th step to producing the correct answer. Then, the total contribution under step-level length control is:
$I_{\text{step}}$ $=\sum_{i}v_{i}\cdot l_{i}^{\text{step}}$ (27)
$=\sum_{i}v_{i}l_{i}-\frac{\alpha}{T}\sum_{i}v_{i}\lambda_{i}l_{i}.$ (28)
Similarly, under global length control, the total contribution is:
$I_{\text{global}}$ $=\sum_{i}v_{i}\cdot l_{i}^{\text{global}}$ (29)
$=\sum_{i}v_{i}l_{i}-\frac{\alpha\lambda}{T}\sum_{i}v_{i}l_{i}.$ (30)
Define $S=\sum_{i}v_{i}l_{i}$. Then:
$I_{\text{step}}-I_{\text{global}}=\frac{\alpha}{T}\left(\lambda S-\sum_{i}v_{i}\lambda_{i}l_{i}\right).$ (31)
Assuming that both methods yield the same total token budget after the update, we have:
$\sum_{i}l_{i}^{\text{step}}=\sum_{i}l_{i}^{\text{global}},$ (32)
and substitute Eq. 25 and 26 into Eq. 32. we can obtain:
$\sum_{i}\lambda_{i}l_{i}=\lambda\sum_{i}l_{i}.$ (33)
In the step-level length control algorithm, $\lambda_{i}$ should be inversely proportional to the contribution density of the step. Thus, we define $\lambda_{i}=1-v_{i}$, yielding:
$I_{\text{step}}-I_{\text{global}}=\sum_{i}v_{i}^{2}l_{i}-\frac{S^{2}}{T}.$ (34)
By the Cauchy–Schwarz inequality:
$\left(\sum_{i}v_{i}^{2}l_{i}\right)\cdot\left(\sum_{i}l_{i}\right)\geq\left(\sum_{i}v_{i}l_{i}\right)^{2}.$ (35)
Therefore, we finally obtain:
$I_{\text{step}}\geq I_{\text{global}}.$ (36)
This result demonstrates that, under the same token budget, step-level length control preserves a higher total contribution toward producing the correct answer compared to global length penalty methods.